# BEADs: Bias Evaluation Across Domains

### v3

@@ Abstract
Recent advancements in large language models (LLMs) have greatly enhanced natural language processing (NLP) applications. Nevertheless, these models often inherit biases from their training data. Despite the availability of various datasets, most are limited to one or two NLP tasks (typically classification or evaluation) and lack comprehensive evaluations across a broader range of NLP tasks. To address this gap, we introduce the Bias Evaluations Across Domains (BEADs) dataset, designed to support a wide array of NLP tasks, including text classification, token classification, bias quantification, and benign language generation. A key focus of this paper is the gold label subset of BEADs, an important portion of the data verified by experts to ensure high reliability. BEADs provides data for both fine-tuning, including classification and language generation tasks, and for evaluating LLMs. Our findings indicate that BEADs effectively identifies numerous biases when fine-tuned on this dataset. It also reduces biases when used for fine-tuning language generation task, while preserving language quality. The results also reveal some prevalent demographic biases in LLMs when BEADs is used for evaluation in demographic task. The benchmarking results highlight the efficacy of fine-tuning LLMs for bias identification and the necessity of comprehensive bias evaluation. We make BEADs publicly available to promote more responsible AI development. The dataset can be accessed at https://huggingface.co/datasets/shainar/BEAD.
@@
Bias , Evaluation , Dataset , Large Language Models
[…]
@@ 1 Introduction
Recent advances in large language models (LLMs) like GPT-4 [1], Llama [2], Mistral [3], Claude [4] have significantly improved NLP tasks such as classification and text generation [5]. However, despite efforts to fine-tune these models for safety [6], they often propagate biases present in their training data [7, 8]. These biases are seen as discriminatory practices related to gender, ethnicity, religion, and other identities in NLP applications. For example, a prompt tested on ChatGPT-4 and Microsoft Copilot multiple runs assumed that a software engineer was male. This example is just one of many observed in NLP applications that demonstrate how biases can manifest in LLMs [9]. This particular example shows a tendency of LLMs to associate technical roles predominantly with males.
@@
Bias is a highly subjective topic, with each discipline defining it differently. Our definition of bias is grounded in social science and computational linguistics [10, 11]. In our work, “bias" covers the language that is discriminatory, potentially offensive, spreads hate speech, or affects sentiments in ways that could be harmful to certain groups. This includes both explicit biases in linguistic nuances and implicit biases conveyed through word choice and tone.
A comprehensive dataset covering various bias aspects (racism, sexism, ageism, among many) is needed for training algorithms to identify and mitigate these biases. Researchers have examined widely-used datasets such as Jigsaw [12, 13], HolisticBias [14], BOLD [15], Stereoset [16], RealToxicityPrompts [17], RedditBias [18] (more given in A) for their potential to identify social biases or toxicity in tasks like bias classification or bias quantification [8, 19].
Despite significant work on biases and NLP datasets, some research gaps remain, which are: (1) Narrow Focus: Existing benchmark datasets often concentrate on only one or two specific NLP tasks, such as bias identification [20, 18] or quantification [16, 14], and somtimes address limited aspects of bias, like gender or race. (2) Limited Labels: High-quality labeled data is needed for training classifiers; however, many datasets rely on a single set of labels. for example, some works [20] is on biases, where as others [12, 17, 21] are on toxicity labels, some works use less precise distant labeling methods (external sources to generate weaker, less accurate labels) [22] for identifying biases. Risk of un-monitored LLM-based Annotations: Although annotation processes have been enhanced by LLMs like OpenAI GPT-3.5/4 [23, 24], however, automating this task without human oversight can result in superficial and further biased outcomes [25]. Lack of Bias Mitigation Data Despite research on the use of prompts with demonstrations for benign text generation [6, 26], there is need for such a dataset for fine-tuning LLMs for benign (unbiased) language generation.
To bridge the identified gaps, we developed Bias Evaluations Across Domains (BEADs), a comprehensive bias evaluation dataset. BEADs is designed to support a wide variety of NLP tasks, including text classification (bias, sentiment, and toxicity labels), token classification (bias entity recognition), and benign language generation (debiasing task) — all of which can be used for training purposes, as well as bias quantification with demographic information (an evaluation task). BEADs provides reliable label sets provided through a combination of GPT-4 annotations and expert verification [23], which is the focus in this paper.
Table 1: Summary of datasets with tasks and annotations by AI or humans. Full comparison in Table LABEL:tab:dataset_comparison.
Dataset Task and Focus Annotation
Toxigen [21] Toxic Classification - Gender, Race, Religion, Disability, Orientation AI & Human
Stereoset Bias Quantification for Gender, Profession, Race, Religion Human
RealToxicity [17] Toxicity Classification (General demographics) Primarily AI
HolisticBias [14] Bias Quantification with demographics: Ability, Age, Body, Characteristics, Culture, Gender, National, Political, Race, Religion, Sexual Orientation, Socio-economic Human
RedditBias [18] Bias Quantification for Religion, Race, Gender, Queerness Human
BOLD [15] Bias Classification for Profession, Gender, Race Human
BEADs Multi-task: Classification (Bias, Toxic, Sentiment), Entity Extraction, Aspect Categorization, Bias Quantification on demo. for Gender, Race, Religion, Mental Health and Language Generation. AI & Human
Given the trend of newer datasets often being in instructional formats for LLM fine-tuning or prompt-tuning [18, 14, 17, 15], BEADs supports a variety of NLP data formats. These include traditional formats like sequence classification in standard dataframes and token classification in BIO tagging / CoNLL format [27], as well as instructional formats such as prompt templates and language generation formats like Alpaca [28]. This versatility facilitates its integration into a wide range of NLP research and applications. BEADs is compared with other bias datasets in Table 1, with detailed comparison in Table LABEL:tab:dataset_comparison.
The main contributions of this work are:
@@
    We developed BEADs, a dataset for bias detection, mitigation and evaluation from social and news media platforms, facilitating various NLP tasks (bias identification, benign language generation and quantification of biases through demographics information).
@@
    We release two versions of the dataset: one initially labeled with GPT-3.5 and enhanced through active learning [29] for scalability (an AI-generated labels approach as in related works [24]), and a high-quality “gold dataset" that combines annotations from GPT-4 that are subjected to experts’ review.
@@
    We fine-tuned open-source smaller language models (such as BERT-like encoder-only models) and LLMs (decoder-only models like Llama/ Mistral) on BEADs for bias classification and language generation task. We also evaluated these models for bias quantification. Our goal is to encourage the integration of BEADs into a broader range of NLP research applications.
Empirical Findings Our findings reveal: (1) Smaller models (encoder only) outperformed LLMs (decoder-only) in the classification tasks, which highlights their efficiency and practicality in real-world applications. (2) However, smaller language models, when evaluated using BEADs portion on demographics, demonstrate more biases in their outputs for certain demographics compared to LLMs. (3) Fine-tuning is more effective than simple prompting with demonstrations. When we fine-tuned LLMs on benign language generation tasks, we observed a significant reduction in bias and toxicity, along with better knowledge retention within the texts, compared to simple prompting.
@@ 2 Literature Review
Named Entity Recognition (NER), though not a widely used method for bias identification, is important methods in identifying biases in NLP [40, 41]. Studies show NER models perform better for certain demographics [42], such as recognizing fewer female names as “PERSON" entities compared to male names in U.S. census data [43]. Some work uses NER to address biases in names and identities (persons, citites etc) by adding learnable adversarial noise to those identities [44]. Counterfactual data augmentation also mitigates gender bias in word embeddings [45, 46]. The NERD framework evaluates linked data entity extractors on semi-structured texts [47]. For some NER tasks, Bi-LSTM with a CRF layer, incorporating pre-trained embeddings like Glove and BERT, and a classification model for sentence splitting is used [48]. The entity debiasing framework for fake news detection models causal relationships among entities, contents, and veracity [49]. Despite various approaches, there is room for dedicated data preparation. We present a dataset identifying bias indicators for use in entity extraction tasks, allowing NLP models to detect biased named entities.
@@
The existing datasets comparisons with BEADs is given in Table LABEL:tab:dataset_comparison.
3 Dataset Construction
This section details the methodologies employed in the creation of the BEADs, from the initial data aggregation to the final benchmarking dataset. For the sake of this work, we define ‘bias’ as language expressing prejudice based on characteristics like gender, race, or religion, often manifesting as toxic language or stereotypes. These definitions are rooted in the related works on societal biases in NLP [59, 60, 16, 61]. We focus on key bias aspects, such as demographic, socio-economic, body image, political , gender, cultural biases, due to their societal significance.
3.1 Data Aggregation
To construct a comprehensive dataset that covers various bias aspects and supports multiple NLP tasks, we compiled data from diverse existing datasets listed in Table 2. We obtained necessary permissions and adhered to their licenses of those existing datasets, as detailed in G. Each dataset was selected for its relevance to specific bias types (political, toxicity, social biases - column 2 of Table 2). We used the texts from these datasets but did not use their original labels, and created new labels for our tasks.
For more aspects to cover, such as occupational, technological, and environmental biases (not covered by the existing datasets), we curated approximately 25,000 news articles via Google RSS from January 1, 2023, to May 30, 2023. The result of combining existing datasets and the newly curated one is a consolidated dataset comprising approximately 3.7 million records, the purpose is to address the limitations of the original sources and provide support for a broad range of NLP tasks.
Table 2: Summary of datasets used in BEADs: All records have been re-labeled using our method.
@@ 3.1 Data Aggregation
Toxic comments [12] Hate Speech and Toxicity 159,571
Jigsaw Unintended Bias Hate Speech and Toxicity [13] 1,800,000
Ageism [22] Health, Ability, and Body Image 887,656
Multi-dimensional news [63] Political 2000
@@ 3.2 Annotation Process
The annotation process is shown in Figure 1, which employs a combination of automated and manual methods to ensure the reliability of the data labels. We developed two versions of the data: (1) initial labeling using GPT3.5 and active learning, and (2) “gold" data with GPT-4-generated labels that were verified by experts. The data has been labeled for various NLP tasks including text classification (bias, sentiment, toxicity), token classification (identifying biased entities, words, phrases), bias aspects (categorization), demographic identification ( for quantifying biases in LLMs), and language generation (creating benign variations of biased texts).
Figure 1: Overview of the BEADs annotation process, which combines automated labeling using GPT-3.5 and active learning for initial data labeling, followed by GPT-4 and manual expert verification to create labels for the gold.
Review Team We assembled a team of 12 volunteer reviewers for our dataset, including four experts and five students from diverse backgrounds in linguistics, psychology, data science, and social sciences. We initially assessed each reviewer’s suitability through a process that evaluated their relevant expertise and commitment to the project. Each team member dedicated specific hours per week to the task. We provided them with clear guidelines on bias, sentiment, and demographic mentions to ensure consistent and high-quality annotations, as detailed in C.
Initial Annotations Initially, we generated annotations using GPT-3.5 with two-shot prompting ( B) to produce preliminary labels on a sample (5̃0k) of our large dataset. We selected this sample size based on text lengths and a mix of texts from each dataset and dimension (focus on bias aspect). These initial labels cover bias, toxicity, and sentiment for text classification. We also generated labels for token classification (e.g., bias words/phrases) and bias aspects (categories like racism, ageism, sexism, cultural, etc.), and identified demographics (race, gender, religion) in the dataset. GPT-3.5 was used initially, since it cost-effective and this annotation method is inspired by related works [23, 65].
To refine and expand these initial labels across the full dataset of 3,674,927 records, we employed an active learning strategy (similar to [24]). During the active learning phase, BERT models were trained on each set of annotated data (for each task) and used to identify instances in the unlabeled dataset. For cases where the ML model exhibited low confidence (we set a threshold of <50%) in its predictions, these instances were presented to an oracle (a human expert, our team managed this with manual workload division) who provided accurate labels for these uncertain cases. The complete details of this initial data labeling is given in B.2.
The main focus in this work is benchmarking on the gold data, the details of which are given below.
Gold Labels For final validation (gold labels), we used GPT-4 in a two-shot demonstration, these propmpts shown in Figure 2 are meant to label multiple tasks. During this phase, we did not use active learning; instead, we relied on human review to ensure annotation accuracy and achieve more reliable labels. The dataset review team oversaw the entire annotation process to ensure high-quality labels across different NLP tasks. They followed detailed guidelines ( C) and conducted regular consensus meetings to resolve discrepancies and confirm label accuracy.
@@
Quality Assurance and Review
To ensure the highest standards of accuracy and consistency for benchmarking task, the labels provided by GPT-4 were thoroughly reviewed by our review team. Discrepancies were resolved through consensus and discussions. To confirm the accuracy and consistency of the gold annotations, we calculated the Cohen Kappa [66] to measure the reliability of the annotator agreements. The annotation results across different tasks are given in Table 3.
Table 3: Annotations Results. Coverage (%) for main data annotation. ‘—’ indicates no data. Gold Labels in ‘k’ for thousands, Demographics as ‘Demo.’, Cohen’s Kappa as ‘Kappa’ for Gold labels.
Task Sub-Task Coverage Gold Labels Kappa
Text Classification Bias, Toxicity, Sentiment 100% (3.67M) 50k 0.67
Token Classification Named Entities 56.39% (2.07M) 27.8k (139.2k Bias tokens) 0.62
Aspects Aspects 75% (2.75M) 50k 0.72
Bias Quantification Token pred. / Stereotypes / Demo. Align. 3% (113.8k) / — / — 29.85k / 352 / 520 0.65 / 0.82 / 0.84
@@
Table 3 presents Kappa values ranging from 0.62 to 0.84 for various tasks , assessing the consistency of our gold annotations. Values below 0.60 prompted corrective actions such as revising annotation guidelines or re-annotating data. On some occasions, a Kappa value of 0.62, close to our threshold, indicated moderate reliability, as observed in related works [67]. Overall, these measures ensured that our annotation process maintained a high standard of consistency and reliability.
In further evaluating GPT-4 as an annotator, we focused on three key qualitative metrics: coverage (capacity to cover a broad range of topics and scenarios within the dataset), consistency (uniformity in labeling across similar data points), and accuracy (correctness of the labels applied). To establish a baseline, a dataset with 500 data points was manually labeled by our review team. The results showed high accuracy at 94% on a set of 500 manually labeled data points, suggesting that GPT-4 labels data correctly the majority of the time. The consistency score of 85% indicates that GPT-4 functions consistently with different kinds of data. Finally, 92% coverage was recorded, indicating that GPT-4 can successfully handle a wide range of subjects and situations in the dataset.
3.3 Gold Data Analysis
Given that benchmarking on gold data is the main focus of this paper, key analysis on the gold dataset is presented in this section. Some preliminary results on the initial labeled dataset analysis is presented in D .

### v4

@@ Abstract
Recent advancements in large language models (LLMs) have greatly enhanced natural language processing (NLP) applications. Nevertheless, these models often inherit biases from their training data. Despite the availability of various datasets for bias detection, most are limited to one or two NLP tasks (typically classification or evaluation) and lack comprehensive evaluations across a broader range of NLP tasks. To address this gap, we introduce the Bias Evaluations Across Domains (BEADs) dataset, designed to support a wide array of NLP tasks, including text classification, token classification, bias quantification, and benign language generation. A key focus of this paper is the gold label dataset that is annotated by GPT4 for scalabilty and verified by experts to ensure high reliability. BEADs provides data for both fine-tuning, including classification and language generation tasks, and for evaluating LLMs. Our findings indicate that BEADs effectively identifies numerous biases when fine-tuned on this dataset. It also reduces biases when used for fine-tuning language generation task, while preserving language quality. The results also reveal some prevalent demographic biases in LLMs when BEADs is used for evaluation in demographic task. We provide the BEADs dataset for detecting biases in various domains, and this dataset is readily usable for responsible AI development and application. The dataset can be accessed at https://huggingface.co/datasets/shainar/BEAD.
@@
bias, large language models, evaluations, promts
@@ 1 Introduction
Large Language Models (LLMs)—such as GPT-4 [1], Llama [2], Mistral [3], and Claude [4], have propelled advances in various NLP tasks, including classification and text generation [5]. Despite ongoing work to prompt-tune or fine-tune these models for safety and fairness [6], they continue to propagate biases inherited from their training data [7, 8]. These biases can manifest as discriminatory behaviors associated with gender, ethnicity, religion, or other identities. For instance, a prompt tested multiple times on ChatGPT-4 and Microsoft Copilot assumed that a software engineer must be male, highlighting how LLMs can implicitly reinforce societal stereotypes [9].
@@
In this work, we define bias from perspectives in social science and computational linguistics [10, 11], encompassing language that is discriminatory, potentially offensive, hateful, or harmful to specific groups. Such biases may appear explicitly in phrasing or implicitly in tone and word choice. The primary objective of this work is to create a dataset designed to detect linguistic biases and serve as a benchmark for evaluating the performance of language models.
To develop robust algorithms for detecting and mitigating bias, comprehensive datasets capturing a wide range of stereotypes related to race, gender, religion, and more are essential. Prior research has introduced datasets such as Jigsaw [12, 13], HolisticBias [14], BOLD [15], Stereoset [16], RealToxicityPrompts [17], and RedditBias [18], which provide important benchmarks for toxicity detection and bias measurement [19]. However, several gaps persist:
First, many benchmark datasets are limited in scope, often focusing on specific tasks (e.g., bias identification or quantification) or narrow bias categories, such as gender or race. Second, few datasets offer high-quality, multifaceted labels, such as bias, toxicity, and sentiment, generated under consistent annotation protocols. Some rely heavily on distant or automated labeling techniques, which can introduce inaccuracies and perpetuate existing biases.
While recent annotation pipelines leverage LLMs for cost-effective scalability [20, 21], the lack of human oversight in automated labeling processes often introduces new biases [22]. Although prompt-based bias mitigation techniques have been explored [6, 23], there remains a need for datasets explicitly designed to support fine-tuning LLMs to generate benign (unbiased) text.
This work seeks to address these challenges by developing a comprehensive dataset that bridges these gaps and advances the field of bias detection and mitigation. We present BEADs (Bias Evaluations Across Domains), a multi-domain dataset designed for a variety of NLP tasks. BEADs includes data for:
  • Classification: Bias, toxicity, and sentiment labels
  • Token Classification: Bias entity recognition
  • Benign Language Generation: Prompt-based debiasing
  • Bias Quantification: Demographic-sensitive evaluation
Our dataset offers reliable label sets derived from a combination of GPT-4 annotations and expert review [20], aiming to balance scalability and accuracy. In line with current trends, BEADs is released in both traditional data formats (e.g., CoNLL/BIO tagging) and instructional formats (e.g., Alpaca [24]) suitable for fine-tuning LLMs. Table 1 presents a snapshot of key bias datasets alongside BEADs. A more detailed comparison can be found in Table  2.
Table 1: Summary of datasets with tasks and annotations (AI or human)
. Dataset Task and Focus Annotation Toxigen [25] Toxic Classification (Gender, Race, Religion, Disability, Orientation) AI & Human Stereoset [16] Bias Quantification (Gender, Profession, Race, Religion) Human RealToxicity [17] Toxicity Classification (General demographics) Primarily AI HolisticBias [14] Bias Quantification (Ability, Age, Body, Culture, Gender, Nationality, Politics, Race, Religion, Sexual Orientation, Socioeconomics) Human RedditBias [18] Bias Quantification (Religion, Race, Gender, Queerness) Human BOLD [15] Bias Classification (Profession, Gender, Race) Human BEADs Multi-task Classification (Bias, Toxicity, Sentiment), Entity Extraction, Demographic-based Bias Quantification, and Debiasing Language Generation AI & Human
1.1 Contributions
The key contributions of this work are:
@@ 1.1 Contributions
    Dataset for Bias Detection and Mitigation: We introduce BEADs, a comprehensive dataset collected from both social and news media. It supports tasks such as bias identification, benign language generation, and demographic-based bias evaluation.
@@
    Combined Annotations: We employ GPT-4 for initial data labeling and then refine annotations via expert review. This ensures scalability and consistency while minimizing automated annotation errors.
@@
    Benchmarking of Fine-Tuned Models: We fine-tune both smaller models (e.g., BERT) and larger LLMs (e.g., Llama/Mistral) on BEADs for classification and generation tasks. Our evaluation also highlights the practical benefits of fine-tuning for bias mitigation (language generation task). We also demonstrates how different model architectures exhibit varying degrees of bias.
1.2 Key Findings
  1. 1.
    Smaller language models vs. LLMs: Smaller, encoder-only (BERT-like) models consistently outperform larger, autogressive decoder-only models (instruct models) in classification, suggesting they are both cost-effective and well-suited for production deployment.
  2. 2.
    Bias Disparities: When generating demographic-sensitive outputs, smaller models display higher rates of bias toward certain demographic groups compared to LLMs, indicating that larger models often incorporate more robust safety guardrails.
  3. 3.
    Fine-Tuning Advantage: Fine-tuning on BEADs on benign language generation task significantly reduces bias and toxicity compared to prompting alone, while preserving domain-specific knowledge and fluency.
Overall, BEADs offers a broad-based resource for the NLP community, supporting research on bias detection, quantification, and mitigation across multiple linguistic domains. We aim for this dataset to catalyze the development of fairer, more responsible language models.
@@ 2 Literature Review
Named Entity Recognition (NER), though not a widely used method for bias identification, is important methods in identifying biases in NLP [36, 37]. Studies show NER models perform better for certain demographics [38], such as recognizing fewer female names as “PERSON” entities compared to male names in U.S. census data [39]. Some work uses NER to address biases in names and identities (persons, citites etc) by adding learnable adversarial noise to those identities [40]. Counterfactual data augmentation also mitigates gender bias in word embeddings [41, 42]. The NERD framework evaluates linked data entity extractors on semi-structured texts [43]. For some NER tasks, Bi-LSTM with a CRF layer, incorporating pre-trained embeddings like Glove and BERT, and a classification model for sentence splitting is used [44]. The entity debiasing framework for fake news detection models causal relationships among entities, contents, and veracity [45]. Despite various approaches, there is room for dedicated data preparation. We present a dataset identifying bias indicators for use in entity extraction tasks, allowing NLP models to detect biased named entities.
@@
The existing datasets comparisons with BEADs is given in Table 2.
Dataset Source Size/Annot./ Open Task Dimension Data Format Year
Racist [55] Twitter 16.9K/H/✓ CLS Hate Speech NLP 2016
HSOL [56] Twitter 24.8K/H/✓ CLS Hate Speech NLP 2017
Winogender [32] Winograd schemas 720/H/✓ CR Gender Bias NLP 2018
WinoBIAS [46] Winograd schemas 3.2K/H/✓ CR Gender Bias NLP 2018
White Supremacy [57] StormFront 9.9K/H/✓ CLS Hate Speech NLP 2018
Toxic Comment [12] Wikipedia 159.6K/H/✓ CLS Toxicity NLP 2018
Age Bias [58] Blogs, Tweets 4.2K blogs, 64.3K comments, 13.8K tweets/H-M/✓ CLS Ageism, Sentiment NLP 2018
Twitter Abusive [59] Twitter 80K/H/NA CLS Hate Speech NLP 2018
Jigsaw [13] Twitter 1.78M/H/✓ CLS Toxicity NLP 2019
Hyperpartisan [60] News publishers 1.3K/H; 754K/✓ CLS Politics NLP 2019
BASIL [61] FOX /NYT /HPO 300/H/✓ CLS Political Bias NLP 2019
SBIC [62] Reddit, Twitter 44.7K/H/✓ CLS Hate Speech NLP 2020
RealToxicity Prompts [17] Webtext corpus 100K/Model/✓ CLS Toxicity Prompts 2020
Stereotype [16] Wikidata 17K/H/✓ BQ Gender, Profession, Race NLP 2020
RedditBIAS [18] Reddit 10.9K/H/✓ LM Eval Religion, Race, Gender NLP 2021
HateMoji [63] Mixed 3.9K/H/✓ CLS Hate Speech NLP 2022
Toxigen [25] GPT-3 274K/Human (9.9K) and Machine (264K)/✓ CLS Implicit Hate Speech NLP, Prompts 2022
HolisticBias [14] Hand-crafted 460K prompts/H/✓ BQ Sociodemographic Bias Prompts 2022
MBIB [64] Various datasets 4.24M/pre-labeled/✓ CLS Multiple Biases NLP 2023
WinoQueer [34] Hybrid 45K/H/✓ BQ Gender, Sexual Orientation NLP 2023
Table 2: Summary of Bias Datasets with Acronyms for Task Column. Acronyms: CLS = Classification, CR = Coreference Resolution, BQ = Bias Quantification, LM Eval = Language Model Evaluation, H = Human annotated
Table 2 provides a concise comparison of bias datasets, and showcasing their diversity in sources, sizes, annotation methods, and bias dimensions. The majority focus on classification tasks (CLS) for identifying hate speech, toxicity, or political bias, with recent datasets incorporating bias quantification (BQ) and sociocultural analysis.
3 Methodology
Figure 1: BEADs Framework
3.1 Task Description
We aim to study and mitigate various types of bias that can appear in text, such as prejudice or unfair assumptions related to gender, race, religion, or other demographic attributes. For the purpose of this work, we follow Baker [65], Dixon et al. [66], Nadeem et al. [16], Navigli et al. [67] in defining ‘bias’ as any language expressing prejudice based on intrinsic characteristics (e.g., gender, race) or cultural/social identifiers (e.g., religion, politics). This definition naturally encompasses manifestations like toxic language or harmful stereotypes. While bias can occur in many forms, we focus on key aspects with substantial societal impact: demographic, socio-economic, body image, political, gender, and cultural biases. The BEADs framework is given in Figure 1, which shows the data collection, task setting, annotation and benchmarking process.
3.2 Dataset Aggregation
This section describes the creation of our BEADs dataset, from initial data collection to the final, consolidated benchmarking resource. Our process involves merging and re-labeling texts from existing bias-related datasets, alongside newly collected data to expand coverage in underrepresented areas.
Source Datasets.
To ensure comprehensive coverage of different bias types, we aggregated text from multiple publicly available datasets, as summarized in Table 3. We obtained the necessary permissions and complied with each dataset’s license (Appendix E). We focused on datasets known for capturing political, toxic, or social biases and used only the raw text from these sources. We did not retain their original labels; instead, we generated new annotations aligned with our own task definitions.
Table 3: Summary of datasets used in BEADs. All records have been newly labeled as part of our methodology.
@@ 3.2 Dataset Aggregation
Toxic Comments [12] Hate Speech & Toxicity 159,571
Jigsaw Unintended Bias [13] Hate Speech & Toxicity 1,800,000
Ageism [58] Health, Ability, & Body Image 887,656
Multi-dimensional News [69] Political 2,000
@@
Final Filtered Dataset Consolidated Bias-Relevant Records 50,000
New Data Collection.
Despite the broad scope of the existing datasets, we found certain bias categories (e.g., occupational, technological, environmental) to be underrepresented. To address these gaps, we scraped approximately 25,000 news articles via Google RSS between January 1 - May 30, 2023. Merging these with the pre-existing data resulted in a consolidated corpus of approximately 3.7 million records.
Filtration Process.
The filtration process was critical in ensuring the quality and relevance of the final dataset. Initially, all records underwent automatic deduplication to remove exact and near-duplicate entries, reducing redundancy. Text quality checks were performed to remove entries with excessive noise, such as malformed text or incomplete sentences. Following a filtration process, the dataset was refined to around 50,000 records. Our aim is for BEADs to serve as a comprehensive resource that addresses limitations in prior datasets and supports a wide array of downstream NLP tasks.
@@ 3.3 Annotation Process
In this work, we utilize automatic annotation through GPT-4. GPT-4, as an annotator, has demonstrated excellent performance in recent studies [20, 70]. However, we include a human-in-the-loop approach to ensure reliability. Figure 1 illustrates our annotation workflow, which combines automated methods (LLM as annotator) with human review to produce high-quality labels. This workflow integrates annotation and quality assurance into a unified process, ensuring both reliability and consistency.
Automated Labeling
We started with automated labeling using GPT-4 in a two-shot demonstration approach (Figure 2). This phase generated provisional labels along with a confidence score of the model on the label, for multiple tasks, including:
  • [leftmargin=*]
  • Text Classification: Identifying bias, toxicity, and sentiment.
  • Token Classification: Marking biased words/phrases and entities.
  • Aspect Categorization: Tagging specific bias types (e.g., political, demographic).
  • Demographic Identification: Labeling data for bias quantification in language models.
  • Language Generation: Generating benign rewrites of biased content.
@@
Human Review and Quality Assurance.
Human review was incorporated as part of the quality assurance process. A team of twelve reviewers from linguistics, psychology, data science, and social science backgrounds ensured diverse perspectives. Each underwent a screening process to confirm expertise and followed standardized guidelines for annotating bias, sentiment, and demographic mentions (Appendix B).
The human review process consisted of several key activities. First, the team reassessed low-confidence automated labels generated by GPT-4 to ensure accuracy. Next, they conducted consensus sessions to reconcile any disagreements among reviewers for collaborative decision-making. Finally, the team re-annotated data where needed to enhance the quality and reliability of the annotations. This iterative process ensured consistency and high-quality gold-standard labels.
To confirm the accuracy and consistency of the gold annotations, we calculated Cohen’s Kappa [71] to measure the reliability of annotator agreements. Cohen’s Kappa was calculated for each task, and the annotation results across different tasks are summarized in Table 4.
Table 4: Annotation Results. Coverage (%) for main data annotations. Gold labels are denoted in ‘k’ (thousands), demographics as ‘Demo.’, and Cohen’s Kappa as ‘Kappa’ for gold labels.
Task Sub-Task Gold Labels Kappa
Text Classification Bias, Toxicity, Sentiment 50k 0.67
Token Classification Named Entities 27.8k (139.2k Bias tokens) 0.62
Aspects Aspects 50k 0.72
Bias Quantification Token pred. / Stereotypes / Demo. Align. 29.85k / 352 / 520 0.65 / 0.82 / 0.84
@@
Table 4 reports Cohen’s Kappa values ranging from 0.62 to 0.84 across tasks, reflecting the consistency of our gold annotations. For tasks with values below 0.60, we took corrective actions, such as revising annotation guidelines or re-annotating data. For instance, a Kappa value of 0.62, though close to the threshold, was deemed acceptable as it aligns with moderate reliability levels reported in related studies [72]. Overall, these measures ensured that our annotation process adhered to high standards of consistency and reliability.
Evaluation of LLM in the Annotator Role
Figure 3: Radar chart showing GPT-4’s annotation performance across Accuracy, Consistency, and Coverage metrics. The chart highlights strong accuracy and coverage, with slightly lower consistency.
To evaluate GPT-4 performance as an annotator, we examined three core metrics through a human evaluation: coverage, consistency, and accuracy. Coverage refers to the model ability to provide annotations across a diverse set of topics and scenarios in the dataset. Consistency measures the uniformity of labeling across similar data points, while accuracy assesses how closely the model’s labels match a gold standard.
To establish a reference for comparison, we manually labeled 500 data points. GPT-4 annotations were then compared against these human-derived gold labels. Figure3 provides a summary of the resulting evaluation. Overall, we observe that GPT-4 demonstrated strong annotating capabilities. Its high accuracy score (94%) indicates substantial alignment with human judgments, while the consistency (85%) and coverage (92%) metrics highlight its reliability across a broad range of data points. Although occasional inconsistencies were noted in particularly nuanced cases, GPT-4 remains a scalable solution for data annotation, with minimal human oversight required to manage edge cases.
Overall, GPT-4 demonstrated strong annotating capabilities. Its high accuracy score (94%) indicates substantial alignment with human judgments, while the consistency (85%) and coverage (92%) metrics highlight its reliability across a broad range of data points. Although occasional inconsistencies were noted in particularly nuanced cases, GPT-4 remains a scalable solution for data annotation, with minimal human oversight required to manage edge cases.
3.4 Dataset Analysis
Text Classification